OpenStudy (anonymous):

lending library has a fixed charge for the first three days and additional charge for day thereafter . sarita paid Rs27for a book kept for 7 days, while susy paid Rs21 for 5 days . find the fixed charge and the charge for extra day.

OpenStudy (anonymous):

What is Rs?

OpenStudy (anonymous):

Rs is rupees

OpenStudy (anonymous):

That's what I thought

OpenStudy (anonymous):

Okay, on to the question, from day 5 to day 7 what is the difference in rupees?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Okay, from that information, what is 6/2

OpenStudy (anonymous):

yup its 3

OpenStudy (anonymous):

From that, it allows you to realize that 3 rupees is the charge after the first 3 days right?

OpenStudy (anonymous):

yes I wanted to know the equation

OpenStudy (anonymous):

Wait, I can't help you then, sorry

OpenStudy (anonymous):

I did it mentally, and I was going to explain my thought process to you

OpenStudy (anonymous):

its okk

OpenStudy (mathstudent55):

This is represented by a linear equation. There is a fixed cost and a daily cost. A linear equation can be written in the slope-intercept form: \(y = mx + b\) where m = slope and b = the y-intercept. We need to find the values of m and b, so you can have the equation. Do you follow so far?

OpenStudy (anonymous):

I did't understood

OpenStudy (mathstudent55):

You stated above you want to know the equation for this problem, but it does not state in the problem that you need the equation. The problem only asks for the daily charge and the fixed charge. I'll continue explaining in the way @Tahamohammed was explaining without an equation. If you still want to find an equation then we'll do that too.

OpenStudy (anonymous):

I know how to do this now with the equation

OpenStudy (anonymous):

Never mind @mathstudent55 you can take over, I'm headed to sleep

OpenStudy (mathstudent55):

Sarita paid Rs27 for 7 days Susy paid Rs21 for 5 days The difference in days between 5 days and 7 days is 2 days (7 - 5 = 2), The difference in total cost between 7 days and 5 days is Rs6 (27 - 21 = 6) Then by dividing Rs6 by 2 days, we get the the daily cost is Rs6/2 = Rs3 per day. This is what @Tahamohammed did above. Since the first 3 days are included in the fixed charge, only the days after the first three days need to be paid with a daily charge. How many days of daily charge did each person use? Sarita had the book for 7 days. 7 - 3 = 4 days. She had the book for 3 days of the fixed charge plus 4 more days of the daily charge. We already know that the daily charge cost Rs3 per day. The 4 days at the daily charge then cost her 4 * Rs3 = Rs12. Since her total charge was Rs27, that means the fixed charge for the first 3 days is Rs27 - Rs 12 = Rs15 The answer is the fixed charge for 3 days is Rs15, and the additional daily charge is Rs3 per day. Now we can check if we are correct using Susy's information. Susy had the book for 5 days. That means 3 days of fixed charge of Rs15 plus 2 days at Rs3 each. Rs15 + 2 * Rs6 = Rs15 + Rs6 = Rs21, which is the given information, so our answer is correct.